### v1

# How much has DESI dark energy evolved since DR1?

Abstract
DESI has reported a dynamical dark energy (DE) signal based on the $w_{0}w_{a}$ model that is in conflict with Hubble tension. Recalling that the combination of DESI DR1 BAO and DR1 full-shape (FS) modeling are consistent with $\Lambda$CDM, in this letter we comment on the status of fluctuations in DR1 BAO documented in Adame et al. (2024a); Ó Colgáin et al. (2024) in the DR2 update. In particular, we note that neither DR1 BAO nor DR2 BAO nor DR2 BAO+CMB confronted to the $w_{0}w_{a}$ model with relaxed model parameter priors confirm late-time accelerated expansion today. Translating DESI BAO constraints into flat $\Lambda$CDM constraints, we observe that the DESI LRG1 constraint remains the most prominent outlier preferring larger $\Omega_{m}$ values, LRG2 switches from smaller to larger $\Omega_{m}$ values relative to Planck-$\Lambda$CDM, and ELG data drive the relatively low $\Omega_{m}$ in the full DR2 BAO. We observe that one cannot restore $w_{0}=-1$ within one $1\sigma$ by removing either LRG1 or LRG2, but LRG2 in DR2, in contrast to LRG1 in DR1, now has a greater bearing on $w_{0}>-1$. We conclude that the BAO has yet to stabilise, but the general trend is towards greater consistency with DESI DR1 FS modeling results, where there may be no dynamical DE signal in DESI data alone.
I Introduction
Recent observations of a statistically significant dynamical dark energy (DE) Adame et al. (2024a, b); Abdul Karim et al. (2025) (see Abbott et al. (2024); Rubin et al. (2023) for earlier claims) based on the CPL model Chevallier and Polarski (2001); Linder (2003) are problematic on a number of fronts. First, local (model-independent) $H_{0}$ determinations are biased to $H_{0}>70$ km/s/Mpc values (see Skara and Perivolaropoulos (2020); Di Valentino et al. (2021); Abdalla et al. (2022); Di Valentino et al. (2025) for reviews). To the extent of our knowledge, it has been appreciated since 2018, perhaps even earlier, that a DE equation of state in the traditional quintessence regime $w_{\textrm{DE}}(z)>-1$ exacerbates the Hubble tension Vagnozzi et al. (2018); Vagnozzi (2020); Alestas et al. (2020); Banerjee et al. (2021); Lee et al. (2022). The result holds in simple $w_{\textrm{DE}}(z)$ parameterizations, e. g. $w$CDM Vagnozzi (2020), $w_{0}w_{a}$ Vagnozzi et al. (2018); Alestas et al. (2020), and more general field theories Banerjee et al. (2021); Lee et al. (2022). In particular, it was observed in Lee et al. (2022) that $w_{0}:=w_{\textrm{DE}}(z=0)>-1$ hinders a resolution to Hubble tension even if $w_{\textrm{DE}}(z)<-1$ at $z>0$. One sees the problem clearly in DESI DR2 results Abdul Karim et al. (2025): in the combination BAO+CMB+SNe, for different SNe samples, respectively Pantheon+ Brout et al. (2022), DES Abbott et al. (2024) and Union3 Rubin et al. (2023), as $w_{0}$ increases, $H_{0}$ decreases. This observation pertains to any combination of datasets with $w_{0}>-1$ and is more general than the CPL model, applying to DE models fitted to DESI data in lieu of CPL, e. g. Lodha et al. (2025a, b).
The second problem is that given the inevitable unknown unknowns in cosmology, one can only trust a result if one sees it independently in distinct datasets. The prototypical example of this is late-time accelerated expansion, which necessitates the presence of DE modeled through $\Lambda$ in the $\Lambda$CDM model, and is seen independently across virtually all observables; not seeing support for $\Lambda$ undermines the observable. In contrast, the DESI dynamical DE claim Adame et al. (2024a, b); Abdul Karim et al. (2025) is only statistically significant when datasets are combined. In defense of combining datasets, it is well recognized that BAO, SNe and CMB only weakly constrain the CPL model on their own. This is in part due to the fact that CPL can be (or should be) viewed as a Taylor expansion in a small parameter $1-a$, which leads to inflated errors relative to other parameterizations Ó Colgáin et al. (2021). Marginalizing over higher order terms also leads to inflated errors in any $w_{0}w_{a}$ model Nesseris et al. (2025).
Even when one restricts attention to independent datasets, a third problem arises: if there is a dynamical DE signal in DESI BAO Adame et al. (2024a); Abdul Karim et al. (2025), when combined with DESI full-shape (FS) modeling Adame et al. (2024b), there may be no dynamical DE signal. To appreciate this, note that Fig. 16 of Adame et al. (2024c) returns consistent constant $\Omega_{m}$ constraints with the $\Lambda$CDM model within $0.7\sigma$. Thus, it is imperative to identify BAO data points that are driving one away from the constant $\Omega_{m}$ behaviour characteristic of the $\Lambda$CDM model Ó Colgáin and Sheikh-Jabbari (2024). Note, in DESI DR2 BAO+CMB, one has a dynamical DE signal at $3.1\sigma$ Abdul Karim et al. (2025). This deviation is expected to be driven by the departures from $\Lambda$CDM behaviour in the relevant late-Universe observable, here BAO. As we shall show, DR1 BAO, DR2 BAO and DR2 BAO+CMB fail to confirm late-time accelerated expansion today.
In Ó Colgáin et al. (2024) the consistency of BAO data was studied by translating $D_{M}(z_{i})/r_{d}$ and $D_{H}(z_{i})/r_{d}$ constraints into direct constraints on the $\Lambda$CDM parameter $\Omega_{m}$ at redshift $z_{i}$. Doing so, it was observed that luminous red galaxy (LRG) data at $z=0.51$ (LRG1) resulted in unexpectedly large $\Omega_{m}$ values relative to Planck Aghanim et al. (2020) at $2.1\sigma$, whereas LRG data at $z=0.706$ (LRG2) led to lower $\Omega_{m}$ values relative to Planck at $1.1\sigma$. In particular, it was easy to argue that LRG1 data was driving the $w_{0}>-1$ signal Adame et al. (2024a); Ó Colgáin et al. (2024). Separately, it was noted that LRG2 distances disagreed with earlier SDSS results Adame et al. (2024a). This allowed one to argue that statistical fluctuations were at work in LRG1 and LRG2 BAO data Adame et al. (2024a); Ó Colgáin et al. (2024) (see also Dinda (2024); Wang (2024); Wang et al. (2024); Chudaykin and Kunz (2024); Liu et al. (2024); Vilardi et al. (2024); Sapone and Nesseris (2024)). These statistical fluctuations disappear when BAO is combined with FS modeling Adame et al. (2024b, c); Ó Colgáin and Sheikh-Jabbari (2024).
Given the obvious conflict with Hubble tension and the risk of statistical fluctuations in LRG BAO data when compared to FS modeling, we revisit earlier analysis. As we show, LRG1 continues to return a $\Lambda$CDM $\Omega_{m}$ value larger than Planck, but at $1.6\sigma$ removed, it is less anomalous. Moreover, LRG2 BAO has flipped from a lower $\Omega_{m}$ value to a higher $\Omega_{m}$ value relative to Planck. This means that while LRG1 drove the $w_{0}>-1$ result in DR1 BAO Ó Colgáin et al. (2024), the $w_{0}>-1$ in DR2 BAO is now driven by both LRG1 and LRG2, but primarily LRG2. On the other hand, whereas the full DESI DR1 BAO dataset preferred a lower $\Omega_{m}$ value relative to Planck driven by LRGs and emission line galaxies (ELGs), in DESI DR2 BAO the lower $\Omega_{m}$ value is driven exclusively by the ELGs. See Ormondroyd et al. (2025) for a recent comparison of DESI DR1 and DR2 BAO. We note that despite the failure to confirm $q_{0}<0$ (without SNe) and the risk of fluctuations, physical modeling of DR2 BAO is underway Brandenberger (2025); Luu et al. (2025); Nakagawa et al. (2025); Paliathanasis (2025); Silva et al. (2025); Wang and Piao (2025); Chaussidon et al. (2025); Kessler et al. (2025); You et al. (2025); Pan et al. (2025); Shlivko et al. (2025).
II Analysis
We begin with a comment on DESI priors $w_{0}\in[-3,1],w_{a}\in[-3,2]$ Adame et al. (2024a); Abdul Karim et al. (2025) when DR2 BAO is confronted with the CPL model. Note, the choice of priors here is arbitrary as there is no theoretical guidance. In contrast, the DES collaboration use more agnostic priors $w_{0}\in[-10,5],w_{a}\in[-20,10]$ Abbott et al. (2024), despite also working with a single observable, namely SNe instead of BAO. By comparing the blue contours in Fig. 1 (DESI priors) to the blue contours in Fig. 2 (DES priors) one sees that the posteriors are skewed with narrower priors. In particular, it is worth noting that the $H_{0}r_{d}$ posterior is positively skewed, while $\Omega_{m}$ and $w_{0}$ posteriors are negatively skewed in Fig. 1. When one defines credible intervals using the most common choices, namely i) equal-tailed intervals and ii) highest density intervals with a mode central value, this skewness manifests itself in terms of larger errors coinciding with the longer tails in the posterior. In contrast, DESI appears to quote the mean as a central value Abdul Karim et al. (2025), $\Omega_{m}=0.352^{+0.041}_{-0.018}$, $w_{0}=-0.48^{+0.35}_{-0.17}$, so the smaller errors coincide with the longer tails. Replacing the mean value with the mode, we find $\Omega_{m}=0.375^{+0.020}_{-0.037}$ and $w_{0}=-0.21^{+0.08}_{-0.44}$, but otherwise we agree with the DESI $68\%$ credible intervals up to small numbers.
Figure 1: CPL model posteriors for DR2 BAO data subject to the DESI priors $w_{0}\in[-3,1]$, $w_{a}\in[-3,2]$ and $w_{0}+w_{a}<0$. The skewness in $(H_{0}r_{d},\Omega_{m},w_{0})$ posteriors comes from the $w_{a}$ bound $w_{a}\geq-3$.
Model+Data $\Omega_{m}$ $w_{0}$ $w_{a}$
CPL DR1 BAO $0.502^{+0.098}_{-0.090}$ $1.03^{+1.0}_{-0.91}$ $-7.3^{+3.2}_{-3.7}$
CPL DR2 BAO $0.385^{+0.046}_{-0.047}$ $-0.19^{+0.44}_{-0.43}$ $-2.7^{+1.5}_{-1.5}$
$z$-exp DR2 BAO $0.372^{+0.036}_{-0.039}$ $-0.47^{+0.29}_{-0.29}$ $-1.19^{+0.66}_{-0.67}$
Table 1: Posteriors for $w_{0}w_{a}$ models and DESI BAO data subject to the priors $w_{0}\in[-10,5]$, $w_{a}\in[-20,10]$ and $w_{0}+w_{a}<0$.
In Fig. 2, having relaxed the priors, we note that the posteriors are visibly more symmetric and Gaussian. Table 1 shows the corresponding parameter constraints. Given the symmetric posteriors, it makes little difference how one defines $68\%$ credible intervals, so we opt for equal-tailed intervals based on percentiles. As Fig. 2 shows, DR2 BAO data confronted to the CPL model is more consistent than DR1 BAO with late-time accelerated expansion today, which requires
$q_{0}=\frac{1}{2}\left[1+3w_{0}(1-\Omega_{m})\right]<0.$ (1)
It is worth noting that $w_{a}$ does not appear in this expression. We examine this requirement [^1] Consistency with the 2011 Physics Nobel Prize, at least at redshift $z=0$, implies a stronger requirement: $q_{0}<0$ at a few $\sigma$ (ideally $5\sigma$ or more). DESI DR2 BAO on its own is not yet of sufficient constraining power to fulfill this requirement within the context of the CPL model. As another comment, our results here can be contrasted with other studies, e.g. Gómez-Valent (2019), where also assuming the CPL model but different data, $q_{0}<0$ is established at $2.5\sigma$. by evaluating the expression for $q_{0}$ on the MCMC chains and examine $q_{0}<0$. See Fig. 3 for the $q_{0}$ posteriors. Converting the MCMC chains into constraints on $q_{0}$, we find that DR1 BAO confronted to the CPL model rules out $q_{0}<0$ at $95.7\%$ confidence level, corresponding to $1.7\sigma$ for a one-sided Gaussian. For DR2 fitted to the CPL model, we find $q_{0}<0$ is ruled out at $76.9\%$ confidence level ($0.7\sigma$ for a one-sided Gaussian). See Wang (2024) for an earlier observation of this tension in DR1. With the improvement in data quality between DR1 and DR2, restrictive priors are no longer required and DR2 BAO shows progress in that $q_{0}$ is less positive.
Combining BAO with external datasets further alleviates this problem. However, even for CMB+DR2 BAO, one can convert $\Omega_{m}=0.353\pm 0.021$, $w_{0}=-0.42\pm 0.21$ Abdul Karim et al. (2025) into $q_{0}=0.09\pm 0.20$, so this result also fails to confirm late-time accelerated expansion today. Only when one combines DR2 BAO with SNe can one confirm late-time accelerated expansion today in a meaningful way. Note, for canonical values of $\Omega_{m}\sim 0.3$, we see that larger values of $w_{0}$ are problematic from equation (1). On the flip side, combining BAO with SNe lowers $w_{0}$ to avoid any contradiction.
Figure 2: Posteriors for $w_{0}w_{a}$ models and DESI BAO data subject to the priors $w_{0}\in[-10,5]$, $w_{a}\in[-20,10]$ and $w_{0}+w_{a}<0$.
As an added check, we replace the CPL model with the analogous $z$-expansion model, $w_{\textrm{DE}}(z)=w_{0}+z\,w_{a}$, which may be viewed as a Taylor expansion in $z$ instead of $1-a=z/(1+z)$ for CPL. One finds a central value more consistent with late-time accelerated expansion today while all the errors shrink accordingly. This reduction in the errors is expected, since as explained in Ó Colgáin et al. (2021), $z$ is a larger expansion parameter than $1-a$ and this allows the data to place stronger constraints on $w_{a}$, which in turn better constrains the remaining parameters. Objectively, the CPL model is a dynamical DE model that is paradoxically less sensitive at lower redshifts. Changing the model from CPL to the $z$-expansion model, we find that $q_{0}<0$ is ruled out at $57.6\%$ confidence level ($0.2\sigma$ for a one-sided Gaussian). See Fig. 3.
Our next task is to identify which data points in DESI DR2 BAO are driving the $w_{0}>-1$ result. One could alternatively focus on the complementary parameter $w_{a}$. We follow the methodology of Ó Colgáin et al. (2024), where for each effective redshift in Table IV of Abdul Karim et al. (2025) with both $D_{M}(z_{i})/r_{d}$ and $D_{H}(z_{i})/r_{d}$ constraints, we construct a $2\times 2$ covariance matrix with the correlation $r$, generate 10,000 $(D_{M}(z_{i})/r_{d},D_{H}(z_{i})/r_{d})$ pairs, and solve the following equation for $\Omega_{m}$ for each pair:
$\frac{D_{M}(z)/r_{d}}{D_{H}(z)/r_{d}}=E(z)\int_{0}^{z}\frac{1}{E(z^{\prime})}\textrm{d}z^{\prime}.$ (2)
This ratio only depends on $\Omega_{m}$ in the $\Lambda$CDM model where $E(z)=\sqrt{1-\Omega_{m}+\Omega_{m}(1+z)^{3}}$. It has been checked that this methodology leads to comparable errors to Markov Chain Monte Carlo (MCMC) Ó Colgáin et al. (2024). The result of this exercise is shown in Table 2 and Fig. 4, where only the lower redshift bright galaxy sample (BGS) is omitted. As remarked in Abdul Karim et al. (2025), this is not a problem as we do not expect strong constraints on $\Omega_{m}$ at lower redshifts. We include the results of LRG3 and ELG1 for completeness in Table 2, but observe that they are not independent from the LRG3+ELG1 entry Abdul Karim et al. (2025).
tracer $z_{\textrm{eff}}$ $\Omega_{m}$
LRG1 $0.510$ $0.467^{+0.11}_{-0.094}$
LRG2 $0.706$ $0.353^{+0.063}_{-0.055}$
LRG3+ELG1 $0.934$ $0.271^{+0.028}_{-0.026}$
ELG2 $1.321$ $0.274^{+0.039}_{-0.033}$
QSO $1.484$ $0.339^{+0.133}_{-0.092}$
Lyman-$\alpha$ QSO $2.330$ $0.304^{+0.037}_{-0.032}$
LRG3 $0.922$ $0.296^{+0.034}_{-0.031}$
ELG1 $0.934$ $0.218^{+0.043}_{-0.038}$
Table 2: Constraints on the $\Lambda$CDM parameter $\Omega_{m}$ from individual tracers.
Figure 3: $q_{0}$ posteriors for the CPL posteriors in Fig. 2.
As remarked earlier, LRG1 data now leads to an $\Omega_{m}$ value that is more consistent with the traditional $\Omega_{m}\sim 0.3$. Relative to the Planck $\Omega_{m}$ value in red in Fig. 4, we see that all constraints intersect the red strip except for LRG1 and LRG3+ELG1. Splitting the LRG3+ELG1 constraint into its components, one sees that this is due to the low $\Omega_{m}$ value ELG1 tracer. Shifting the red horizontal strip downwards to the location of the green strip corresponding to the DESI DR2 BAO $\Omega_{m}$ constraint for the full sample, one can see that this provides visually a better fit to all constraints, except the LRG1 constraint, which returns a larger $\Omega_{m}$ value at $1.8\sigma$ [^2] Removing the LRG1 data we have $\Omega_{m}=0.293\pm 0.009$ in the $\Lambda$CDM model, so the discrepancy between LRG1 constraint in Table 2 and the remaining data is also $1.8\sigma$. This makes LRG1 the most prominent outlier once again. It should be noted that in contrast to DR1, where LRG2 data was contributing to the lower $\Omega_{m}$ in the full BAO sample, we can now confirm that this is driven exclusively by ELG data. It is also worth noting that the LRG3+ELG1 constraint has exhibited the smallest shift in $\Omega_{m}$ value between DR1 and DR2.
We now ask again the question posed in Ó Colgáin et al. (2024): is it possible to remove a single constraint, for example LRG1, and recover $w_{0}=-1$ (or $w_{a}=0$) within $1\sigma$? In Table 3 we document the effect of removing LRG1 and LRG2 data, which are the most obvious points that could be driving the $w_{0}>-1$ result in the full sample. Interestingly, we find that the removal of LRG1 pushes one further into a regime where there is no accelerated expansion today. In contrast, we find that removing LRG2 brings the CPL parameters to within $1.3\sigma$ of $\Lambda$ ($w_{0}=-1,w_{a}=0$). What this implies is that the $w_{0}>-1$ deviation from $\Lambda$CDM in DR2 BAO is driven by LRG2 data, whereas in DR1 BAO, this was driven by LRG1 data. The reader should note that LRG1 appears to moderate the large $w_{0}$ value attributable to LRG2 data. Indeed, with LRG1 data removed, so that only LRG2 is driving the $w_{0}>-1$ deviation, we find that $q_{0}<0$ is disfavoured at $87.3\%$ confidence level ($1.1\sigma$ for a one-sided Gaussian).
In summary, our analysis exposes a $q_{0}$ sign problem and evident fluctuations in DESI DR2 BAO. As is clear from Fig. 4, LRG3+ELG1 shows excellent agreement between DR1 and DR2, while ELG2 and Lyman-$\alpha$ QSO also show good agreement. On the other hand, LRG1 has become more consistent with Planck. The most interesting difference between DR1 and DR2 is the shift in LRG2, which means it goes from contributing to the lower $\Omega_{m}$ preferred by the full sample relative to Planck, to the tracer that is most sharply driving the dynamical DE signal. In Fig. 5 we remind the reader that despite the potential for statistical fluctuations in DESI BAO data, when DESI DR1 BAO is combined with FS modeling, there are no obvious outliers and all constraints intersect the $\Omega_{m}$ value for the full BAO+FS dataset Adame et al. (2024b, c). As a result, there is no signal of dynamical DE in DESI DR1 data alone. That being said, DESI DR1 data prefers a value for the $\Lambda$CDM parameter $\Omega_{m}$ that is $1.6\sigma$ lower than the Planck value. Admittedly, this constant shift in $\Omega_{m}$ challenges concordance, but claims that $\Omega_{m}$ is not a constant in $\Lambda$CDM cosmology, and is in fact redshift dependent, go back to 2022 Ó Colgáin et al. (2022, 2024). The main point here is that a non-constant $\Omega_{m}$ $\Lambda$CDM parameter, while pointing to model breakdown, does not immediately imply a dynamical DE sector.
Data $\Omega_{m}$ $w_{0}$ $w_{a}$
no LRG1 $0.418^{+0.066}_{-0.062}$ $0.20^{+0.68}_{-0.63}$ $-4.1^{+2.1}_{-2.3}$
no LRG2 $0.363^{+0.049}_{-0.054}$ $-0.42^{+0.47}_{-0.46}$ $-2.0^{+1.7}_{-1.6}$
Table 3: Constraints on the CPL model from the full sample with LRG1 and LRG2 data removed.
Figure 4: Differences in the $\Lambda$CDM $\Omega_{m}$ constraints from individual tracers between DR1 in faded blue and DR2 in blue. The red and green bands denote Planck and full DESI DR2 BAO sample constraints on $\Omega_{m}$. In magenta we separate LRG3 and ELG1 constraints.
Figure 5: DESI DR1 BAO constraints on the $\Lambda$CDM parameter $\Omega_{m}$ in faded blue relative to the DESI DR1 BAO+FS modeling constraints in blue. The green strip denotes the constraint on the full sample from BAO+FS and it can be confirmed that all constraints show excellent agreement.
III Discussion
Any dynamical DE signal with $w_{0}>-1$ cannot be the final word on a replacement for the $\Lambda$CDM model because the $w_{0}-H_{0}$ anti-correlation is problematic in the face of $H_{0}>70$ km/s/Mpc determinations Vagnozzi et al. (2018); Vagnozzi (2020); Alestas et al. (2020); Banerjee et al. (2021); Lee et al. (2022). A key point is that larger than expected local $H_{0}$ values are observed across multiple observables. As a result, one infers that there must be something wrong with the DESI dynamical DE claim Adame et al. (2024a, b); Abdul Karim et al. (2025), at least in its current form. It is true that the CPL model fits DESI+CMB+SNe and DESI+CMB datasets better than $\Lambda$CDM in a statistically significant manner, but if it contradicts Hubble tension, or worse, fails to confirm late-time accelerated expansion today, one should bin the idea. That being said, it should be borne in mind that Hubble tension is an expected harbinger new physics beyond $\Lambda$CDM.
In this letter we looked at the improvements in DESI BAO data between DR1 Adame et al. (2024a) and DR2 Abdul Karim et al. (2025). When the data is confronted to the CPL model, we noted that the narrow $w_{a}\in[-3,2]$ priors employed by the DESI collaboration cut off the $w_{a}$ posterior, giving rise to skewed posteriors that become more symmetric once the priors are relaxed. In agreement with earlier work Wang (2024), with relaxed priors we confirmed that DESI DR1 BAO is inconsistent with late-time accelerated expansion today ($q_{0}<0$). In the upgrade to DESI DR2 BAO, we note that $q_{0}<0$ cannot be confirmed, even when one combines DR2 BAO with CMB. Thus, we have a $3.1\sigma$ deviation from $\Lambda$CDM Abdul Karim et al. (2025), yet cannot confirm $q_{0}<0$.
Given that DESI DR1 BAO data are prone to fluctuations Adame et al. (2024a); Ó Colgáin et al. (2024), and the dynamical DE signal may have hinged on an isolated LRG1 tracer Ó Colgáin et al. (2024), we revisited earlier analysis that translates $(D_{M}(z_{i})/r_{d},D_{H}(z_{i})/r_{d})$ pairs at redshift $z_{i}$ into constraints on the $\Lambda$CDM parameter $\Omega_{m}$. This is an important exercise as the statistically significant deviation from $\Lambda$CDM reported in DR2 BAO+CMB at $3.1\sigma$ Abdul Karim et al. (2025) but $q_{0}>0$ must be due to BAO data.
We observe that LRG1, which was the most prominent outlier in DR1 BAO Ó Colgáin et al. (2024), is now consistent with a lower $\Omega_{m}$ value in the $\Lambda$CDM model. It is still the most prominent outlier at $1.8\sigma$ removed from the rest of the dataset. Moreover, we find that LRG2 returns an $\Omega_{m}$ value larger than Planck in DR2 in contrast to the smaller value in DR1. We now confirm that ELG data is solely responsible for the lower $\Omega_{m}$ relative to Planck in the full DR2 BAO sample. Finally, we show that LRG2 and not LRG1 is now most responsible for the $w_{0}>-1$ dynamical DE signal.
Evidently, fluctuations still persist in DESI BAO data and we have yet to see convergence in constraints for both LRG1 and LRG2, although high redshift tracers show good to excellent agreement between DR1 and DR2. It will be interesting to see if any fluctuations remain when DR2 BAO is combined with FS modeling, as there is no dynamical DE signal when DR1 BAO is combined with DR1 FS modeling Adame et al. (2024c).
It is important to note that a dynamical DE signal can have two origins. In BAO+CMB+SNe combinations, differences in the $\Lambda$CDM parameter $\Omega_{m}$ between datasets at different effective redshifts can manifest as a dynamical DE signal even if there is no dynamical DE signal in BAO and SNe independently. See Ó Colgáin and Sheikh-Jabbari (2024) for consistency checks of this possibility. The second possibility is that there is a genuine dynamical DE signal in these independent datasets, one currently seen in DES SNe Abbott et al. (2024), DESI BAO Adame et al. (2024a); Abdul Karim et al. (2025), but importantly not DESI FS modeling Adame et al. (2024b, c). The community needs to separate these two possibilities and study them independently. Nevertheless, no matter how one looks at it, $w_{0}>-1$ has a Hubble tension problem Vagnozzi et al. (2018); Vagnozzi (2020); Alestas et al. (2020); Banerjee et al. (2021); Lee et al. (2022), which risks making any discussion moot.
Acknowledgements
We would like to thank Eleonora Di Valentino and Sunny Vagnozzi for correspondence on credible intervals. This article/publication is based upon work from COST Action CA21136 – “Addressing observational tensions in cosmology with systematics and fundamental physics (CosmoVerse)”, supported by COST (European Cooperation in Science and Technology).